OpenStudy (anonymous):

use dimensional analysis to find a formula for the time it takes the marble to return to the surface of the earth

OpenStudy (anonymous):

What marble? What's the setup?

OpenStudy (anonymous):

that's all it gave me

OpenStudy (anonymous):

If that was it, it would say "a" marble, not "the" marble.... was it thrown up? Just dropped?

OpenStudy (anonymous):

oh it does say "a" marble sorry, it says it is shoot up

OpenStudy (anonymous):

Okay so what do you think the time might depend on?

OpenStudy (anonymous):

What variables, I mean?

OpenStudy (anonymous):

seconds

OpenStudy (anonymous):

No, that's the unit that the answer will be in, but what variables would the time of flight depend on? Like maybe the mass of the marble, or the time of day you shoot it up.... :)

OpenStudy (anonymous):

well it does say but i could guess gravity?

OpenStudy (anonymous):

That would be a good guess, anything else?

OpenStudy (anonymous):

air resistance but i think this case we would neglect that

OpenStudy (anonymous):

True... one more thing, it's staring you in the face... :)

OpenStudy (anonymous):

the obvious ones are always so hard to think of... maybe the velocity that it is shot?

OpenStudy (anonymous):

There you go :) So let's say that \[T \propto g^x\cdot v^y \]

OpenStudy (anonymous):

What are the units on the left?

OpenStudy (anonymous):

time in seconds

OpenStudy (anonymous):

Great, now what about on the right? for example, \[g^\alpha\] has units of \[m^\alpha\cdot s^{-2\alpha} \]

OpenStudy (anonymous):

i'm so sorry for doing that, but please, can someone help me?

OpenStudy (anonymous):

mass, m/s and m/s^2 i think

OpenStudy (anonymous):

m is meters, a unit of length. The way to do it goes like this: I showed you the units that we get from the g. Including the velocity, we get this: \[ \text{units of } g^x v^y = \left(\frac{m}{s^2}\right)^x \cdot \left( \frac{m}{s}\right)^y = m^{x+y}{s^{-(2x+y)}}\]

OpenStudy (anonymous):

This must equal seconds to the power 1, so we can solve these two equations and figure out x and y. Since we have no unit of length, \[x+ y = 0 \rightarrow x = -y\] Plugging that in to the second equation and setting it equal to +1, \[-(2(-y) + y) = -(-y) = y = 1 \] So y = 1, x = -1, and apparently \[T \propto \frac{v}{g} \]

OpenStudy (anonymous):

As it turns out, the real answer is \[T = 2\frac{v}{g} \] but dimension analysis always gives us proportional relationship that doesn't take into account dimensionless factors like 2.

OpenStudy (anonymous):

okay makes sense thanks a lot, i know dimension analysis is supposed to make everything easier but it never clicked for me anyway thanks again

OpenStudy (anonymous):

No problem. It doesn't necessarily make problems easier, per se, but it provides a good way to check if your formula is correct, and it is a comparatively fast way to squeeze some information out of the problem without actually solving it.